# SafeGen: Mitigating Sexually Explicit Content Generation in Text-to-Image Models

### v2

@@ 6.1.1. Nudity Content Reduction
Overall effectiveness. Table 1 shows that SafeGen outperforms the baselines by achieving the highest average NRR of 95.6% across across all adversarial prompts. The baselines exhibit a range of NRR values, with the lowest being 8.5% (SLD (Weak)) to 72.1% (ESD), averaging at 49.8%. In addition, a visual comparison between SafeGen and these baselines provided by Figure 15 further demonstrates the effectiveness of SafeGen.
We have three observations. First, external methods, on average, successfully remove 60.2% of nude content. However, due to the limitations of filters used in training data censorship or inference-stage filtering, particularly those involving less obvious content, out-of-distribution explicit content, and perturbations such as SneakyPrompt-P with pseudo words, may evade detection. Second, text-dependent mitigation can remove 70.7% nude content on average if we only consider those methods with the highest safety level, i.e., ESD and SLD (Max). While ESD manipulates the model weights to erase predefined textual unsafe concepts, it may not account for all variations of such content or new content that evolve over time (e.g., porn stars’ names), leading to less effectiveness in nudity removal. The difficult-to-enumerate challenge also limits the performance of SLD. Third, it is worthwhile to mention that SafeGen archives an impressive performance on the NSFW-56k dataset, i.e., 99.1% NRR. In contrast, the other baselines show different degrees of effectiveness, e.g., from 5.6% (SLD (Weak)) to 70.0% (safety filter). These outcomes suggest that the NSFW-56K dataset may serve as a challenging benchmark for future works in this domain.
Different nude body parts. Figure 10 shows the results of SafeGen and baseline methods in reducing the generation of various exposed body parts, e.g., M-Breasts or F-Breasts, on the NSFW-56k dataset, where ‘M’ stands for male and ‘F’ stands for female. SafeGen achieves a 99.1% NRR for total exposed body parts, while the others are less effective on some body parts. For example, SLD (Strong) exhibits a 22.2% NRR for buttocks, and SD-V2.1 has a 26.0% NRR for belly, which indicates their limitation on undefined or unseen NSFW concepts. Moreover, a -16.9% NRR on buttocks caused by SLD (Weak) suggests some safe measures can unintentionally steer the denoising diffusion process towards unsafe regions. Due to the page limitation, we display the removal results on other three datasets in Appendix E (Figure 16).
@@ 6.3. RQ3: Performance Combined with Baselines
Table 4 shows the results of the performance of SafeGen when combined with baselines. We evaluate the combination with text-dependent baselines, i.e., ESD and different variants of SLD on both nudity mitigation and benign preservation. We skip the safety filter baseline since it has no impact on benign generation. We employ SneakyPrompt-N for testing nudity mitigation and COCO-25k for testing benign preservation. In each dataset, we randomly select 200 prompts, and then generate three images per prompt using different random seeds. Our findings reveal that SafeGen, with only self-attention layers adjustment, alone outperforms all baselines in terms of nudity removal with average 21.7% NRR improvement, while retaining high-fidelity benign generation with comparable CLIP score. In addition, the integration with other text-dependent techniques demonstrate SafeGen significantly aids baselines in reducing sexually explicit content generation, realizing a remarkable 27.8% NRR enhancement.
From the perspective of nudity content mitigation, SafeGen + SLD (Max) achieves the highest NRR at 97.8% and the lowest CLIP score at 16.75, indicating its effectiveness in mitigating exposed body parts generation and deviating the resulting images from adversarial prompts. On the other hand, from the perspective of benign generation mitigation, SafeGen + SLD (Weak) has the lowest LPIPS score at 0.787 and the highest CLIP score at 24.33, which suggests it preserves the visual fidelity of benign images well.
@@ 6.4.1. Different Hyperparameters of Loss Weights
Table 5 presents the performance of SafeGen on both nudity removal and benign preservation by varying the loss weights $\lambda_{m}$ and $\lambda_{p}$. We observe an overall upward trend of NRRs and downward trend of CLIP scores by gradually increasing $\lambda_{m}$ and decreasing $\lambda_{p}$, denoting a larger $\lambda_{m}$ and a smaller $\lambda_{p}$ benefits nudity content removal and suppresses the adversarial text-to-image alignment. The average NRR and CLIP score under different $\lambda_{m,p}$ combinations are up to 95.9% and down to 17.57, with negligible variance, suggesting SafeGen can well mitigate nudity concepts with a wide parameter space. Prompted by benign texts from the COCO-25k dataset, results demonstrate a lower $\lambda_{m}$ and higher $\lambda_{p}$ can ensure that SafeGen yields high-fidelity images. The optimal LPIPS score is down to 0.789 and the best CLIP score reaches 24.60, even slightly surpasses the original SD’s performance with 24.56. This experiment guides us in selecting the optimal hyperparameter combination, i.e., $\lambda_{m}$: 0.1 and $\lambda_{p}$: 0.9, that excels in defense efficacy while preserving the fidelity of benign image generation from a systematic performance perspective.

### v3

@@ 6.1.1. Nudity Content Reduction
Overall effectiveness. Table 1 shows that SafeGen outperforms the baselines by achieving the highest average NRR of 97.1% across across all adversarial prompts. The baselines exhibit a range of NRR values, with the lowest being 5.2% (SLD (Weak)) to 85.3% (ESD), averaging at 55.5%. In addition, a visual comparison between SafeGen and these baselines provided by Figure 15 further demonstrates the effectiveness of SafeGen.
We have three observations. First, external methods, on average, successfully remove 65.4% of nude content. However, due to the limitations of filters used in training data censorship or inference-stage filtering, particularly those involving less obvious content, out-of-distribution explicit content, and perturbations such as SneakyPrompt-P with pseudo words, may evade detection. Second, text-dependent mitigation can remove 78.3% nude content on average if we only consider those methods with the highest safety level, i.e., ESD and SLD (Max). While ESD manipulates the model weights to erase predefined textual unsafe concepts, it may not account for all variations of such content or new content that evolve over time (e.g., porn stars’ names), leading to less effectiveness in nudity removal. The difficult-to-enumerate challenge also limits the performance of SLD. Third, it is worthwhile to mention that SafeGen archives an impressive performance on the NSFW-56k dataset, i.e., 99.4% NRR. In contrast, the other baselines show different degrees of effectiveness, e.g., from 5.2% (SLD (Weak)) to 74.4% (ESD). These outcomes suggest that the NSFW-56K dataset may serve as a challenging benchmark for future works in this domain.
Different nude body parts. Figure 10 shows the results of SafeGen and baseline methods in reducing the generation of various exposed body parts, e.g., M-Breasts or F-Breasts, on the NSFW-56k dataset, where ‘M’ stands for male and ‘F’ stands for female. SafeGen achieves a 99.4% NRR for total exposed body parts, while the others are less effective on some body parts. For example, SLD (Strong) exhibits a 22.2% NRR for buttocks, and SD-V2.1 has a 26.0% NRR for belly, which indicates their limitation on undefined or unseen NSFW concepts. Moreover, a -16.9% NRR on buttocks caused by SLD (Weak) suggests some safe measures can unintentionally steer the denoising diffusion process towards unsafe regions. Due to the page limitation, we display the removal results on other three datasets in Appendix E (Figure 16).
@@ 6.3. RQ3: Performance Combined with Baselines
Table 4 shows the results of the performance of SafeGen when combined with baselines. We evaluate the combination with text-dependent baselines, i.e., ESD and different variants of SLD on both nudity mitigation and benign preservation. We skip the safety filter baseline since it has no impact on benign generation. We employ SneakyPrompt-N for testing nudity mitigation and COCO-25k for testing benign preservation. In each dataset, we randomly select 200 prompts, and then generate three images per prompt using different random seeds. Our findings reveal that SafeGen, with only self-attention layers adjustment, alone outperforms all baselines in terms of nudity removal with average 18.8% NRR improvement, while retaining high-fidelity benign generation with comparable CLIP score. In addition, the integration with other text-dependent techniques demonstrate SafeGen significantly aids baselines in reducing sexually explicit content generation, realizing a remarkable 24.2% NRR enhancement.
From the perspective of nudity content mitigation, SafeGen + SLD (Max) achieves the highest NRR at 98.2% and the lowest CLIP score at 16.75, indicating its effectiveness in mitigating exposed body parts generation and deviating the resulting images from adversarial prompts. On the other hand, from the perspective of benign generation mitigation, SafeGen + SLD (Weak) has the lowest LPIPS score at 0.787 and the highest CLIP score at 24.33, which suggests it preserves the visual fidelity of benign images well.
@@ 6.4.1. Different Hyperparameters of Loss Weights
Table 5 presents the performance of SafeGen on both nudity removal and benign preservation by varying the loss weights $\lambda_{m}$ and $\lambda_{p}$. We observe an overall upward trend of NRRs and downward trend of CLIP scores by gradually increasing $\lambda_{m}$ and decreasing $\lambda_{p}$, denoting a larger $\lambda_{m}$ and a smaller $\lambda_{p}$ benefits nudity content removal and suppresses the adversarial text-to-image alignment. The average NRR and CLIP score under different $\lambda_{m,p}$ combinations are up to 97.6% and down to 17.57, with negligible variance, suggesting SafeGen can well mitigate nudity concepts with a wide parameter space. Prompted by benign texts from the COCO-25k dataset, results demonstrate a lower $\lambda_{m}$ and higher $\lambda_{p}$ can ensure that SafeGen yields high-fidelity images. The optimal LPIPS score is down to 0.789 and the best CLIP score reaches 24.60, even slightly surpasses the original SD’s performance with 24.56. This experiment guides us in selecting the optimal hyperparameter combination, i.e., $\lambda_{m}$: 0.1 and $\lambda_{p}$: 0.9, that excels in defense efficacy while preserving the fidelity of benign image generation from a systematic performance perspective.